OpenStudy (anonymous):

can someone show me how to factor 6x^2 - 3xy - 3y^2

OpenStudy (tkhunny):

Why do you believe it CAN be factored?

OpenStudy (anonymous):

because my book has the answer in the answer key

OpenStudy (tkhunny):

That's a pretty good reason, but hardly scriptural. You can narrow it down. 3 is prime. Your answer MUST look like this: (__ x + y)(__ x - 3y) You just have to find the right factors of 6 to make it work. 1, 6 6, 1 2, 3 3, 2 There you have it. Only four things to try.

OpenStudy (anonymous):

hmm the answer is 3(2x + y)(x - y)

OpenStudy (tkhunny):

Try 2, 3 from my list.

OpenStudy (anonymous):

well that would make (2x + y) (3x - 3y) thats not the answer though?

OpenStudy (tkhunny):

Have you considered the Distributive Property? (3x-3y) = 3(x-y)

OpenStudy (anonymous):

sorry i dont understand that would still make (2x + y) 3(x - y) no?

OpenStudy (tkhunny):

Have you considered the commutative property? (2x+y)3(x-y) = 3(2x+y)(x-y) These are important properties. You are expected to have them in your bag of tricks. Also, I failed to notice that all the coefficients were divisible by 3. This would have made our lives easier. 6x^2 - 3xy - 3y^2 = 3(2x^2 - xy - y^2) = One of these 3(2x + y)(x - y) or 3(2x - y)(x + y)

OpenStudy (anonymous):